RC Meeting, Wednesday, October 20 , 2021

From HallCWiki
Jump to: navigation, search


RC Meeting Information

RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920

Accelerator Status

  1. Accelerator went down after LCW issue in the SAB last night. Water conductivity remained too high to restore beam until this morning. Photos: https://logbooks.jlab.org/entry/3928348
  2. Accelerator maintenance day underway
    1. Work in tunnel until mid-afternoon
    2. Lock up machine, RF recovery (several hours)
    3. Beam restoration
    4. Beam to halls by midnight (if all goes well)
  3. Another pass change on Friday (Hall A to 4 pass) -> maybe this will change?

Hall C Status

  1. So far we have collected about 3868 pions (as of 7:30 am) in this configuration
  2. Hall C is in Controlled Access

Current Run Plan

  1. Upon beam restoration, execute the abbreviated beam restoration procedure outlined here: https://logbooks.jlab.org/entry/3928364
  2. Continue data taking until early Nov

Current Issues

  1. Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
  2. HB LHe has been stable since Saturday owl -> some valve adjustments on Monday
  3. Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
  4. Still running with no UPS on the SHMS cryo controls - Steve L. will put in PR for batteries and UPS. ( PENDING )
  5. HCCAM02 not functioning (lift required to replace).

Special Requests

Useful Links